Prev Next

SW&IT Meeting Minutes 2020-12-07

Attendees: Carl, David, Igor, Matthieu, Stephen, Trey, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • ATT SW complete 95%
    • Argos cal-unit in TCS mini-review
    • sci-arch meeting
  • This week
    • ATT SW / ATT PM
    • leadership workshop Tue/Wed 8am-12pm
    • Carl's responsibilities within the group
    • iLocater fiber discussions?

DE

  • Office visit
    • HD swap
    • UPS battery tray collect
    • Added UPS to Rack 3, splitting load between 2 units
    • KVM dongle; decided not to proceed
    • Check CH office for returned equipment
  • User account work
    • Google will force a reset on your password if you use bad practice and have a reused password compromised. Neat. Don't reuse passwords.
  • UA IT training/disappointing
  • Cisco training/much more satisfying
This week:
  • Rebuild UPS batteries
    • Batteries Plus will accept the old SLA batts for free; will drop them off
  • Cisco
  • User cleanup/CH/SB?

YZ

  • Previous:
    • Some LBC work (DMS updates, LBC/IA study, RB_Science updates etc).
    • Some TMS work (Utility release, software updates, wiki edits, and meeting).
    • Some AO-Cravat/OSCO work (mainly dash framework study).
  • This week:
    • Some misc LBC and MODS work to continue.
    • Some TMS work and possible software support.
    • Mainly focus on AO-Cravat/OSCO work.
    • UA information security awareness certification due soon?

SH

  • Observer support
  • grouper stuff with UITS
  • Leroy PEPSI and obs computers
  • SSL for various entities
  • submit final MAC addresses
  • mountain work for second robs and vm2 HV.
  • cleanup of info pages and things
  • CameraLink quotes
  • vm2 HV bad USB?

next:

  • CameraLink purchase
  • Help install ovirt manager
  • vm2 HV
  • Fedora installs on PEPSI recon units

XZ

  • Last week
    • Setup the lab enviroment for flashing the DSP firmware with Juan Pablo and Guido
    • Revert and test the implementation on AOS side for elevation limit change
    • ATT SW
    • Completed UA harassment training
    • ATT AdSec
    • AO operation support
  • This week
    • ATT AdSec
    • ATT SW
    • Change ASM Elevation Limits & Configuration demonstration and dicussion
    • Implementation on ao-suervisor side for elevation limit change
    • AO operation support

IS

Last week:
  • Harassment training
  • LUCI
    • RPMs generation.
    • Continue Geirs script test Barry's night log.

This week:
  • Harassment training
  • LUCI
    • RPM
    • IT 8231 GEIRS timeout
    • IT 8220
    • split current luci motor privilege into two to allow more control on what people can and can't do.
    • Look into reading WebIO status and passing it to the database/telemetry.
    • Add more debugging information in Journalizer DB storing logic. Trying to figure out why sometimes the mask config file doesn't upload.
  • PIT
    • pit admin tool

PSG

  • Last week
    • Talk with Gustavo to define parameters for swing arms and white light source control.
    • Begin coding software to control both.
  • This week
    • Continue coding.
Topic revision: r7 - 08 Dec 2020, PaulGrenz
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback